OpenStudy (anonymous):

Will someone please tell me what you do, step by step to solve a quadratic equation. I am having trouble specifically with the discriminant, and with completing the equation for the solutions. I have the formula memorized, so no problem there.

OpenStudy (anonymous):

I'm not sure what you mean. If you have the formula, at which point do you get stuck? Like, where do you get stuck trying to find the zeroes of: \[x ^{2} - x + 1\]

OpenStudy (anonymous):

Here is an example: \[x^2 -4.4x +2.8=0\] i can get all the way to finding the discriminant, but i have no idea how to find the solution after that.

OpenStudy (anonymous):

i am sitting at \[x= 4.4 + \sqrt{43.008} \] and \[x= 4.4-\sqrt{43.008}\] both divided by two

OpenStudy (anonymous):

You made a mistake under the root. The formula at that point is \[\sqrt{b ^{2}-4ac}\]

OpenStudy (anonymous):

b in this case being -4.4, a =1 and c = 2.8

OpenStudy (anonymous):

which comes out to be 8.16

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i still do not know what to do with it after this point

OpenStudy (anonymous):

There is not much you can do, without a calculator, I don't think they want you to solve the root of 8.16 by hand

OpenStudy (anonymous):

i see the error by the way, thank you for pointing that out

OpenStudy (anonymous):

No problem, the results of x at y=0 are then x=\[4.4+\sqrt{8.16}\over2\] and x=\[4.4-\sqrt{8.16}\over2\]

OpenStudy (anonymous):

it doesn't simplify further?

OpenStudy (anonymous):

you can make that prettier when you divide 4.4 by 2 and take 1/2 into the root which then becoms 1/4 and divide 8.16 by four. But that's it. Unless you are allowed to use a calculator and take the root of 8.16

OpenStudy (anonymous):

Need the working?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

it needs to be simplified

OpenStudy (anonymous):

Ok, so you can write the fracture into: \[4.4\div2 + \sqrt{8.16}\div2\]

OpenStudy (anonymous):

so: \[2.2 + \sqrt{4.08}\] and \[2.2-\sqrt{4.08}\] both being divided by 2

OpenStudy (anonymous):

Then you can square 2 which gives four and put in under a root (then it would be two again, but you leave it at that. Then you when its division or multiplication, you are allowed to divide first, then take the root. This gives: \[2.2 + \sqrt{8.16 \over 4}\]

OpenStudy (anonymous):

Yes, you did this right!

OpenStudy (anonymous):

Oh, no sorry, too quick! You forgot one step. You cannot change the value of 2 in this equation, so under the root it becomes four!! That is important. Do you understand?

OpenStudy (anonymous):

the answer is 3.63, 0.77 but how in the world did they get that?

OpenStudy (anonymous):

Well, with a calculator you can just type in root of 2.04, and add it to 2.2 for the first x and subtract it to get the second.

OpenStudy (anonymous):

This gives you the numbers in the answers.

OpenStudy (anonymous):

the root of 2.04?

OpenStudy (anonymous):

Yes, 8.16 divided by four gives 2.04.

OpenStudy (anonymous):

ok, how did you know to divide it by 4?

OpenStudy (anonymous):

Ok, so in the equation it says \[\sqrt{8.16}\over2 \] are you with me?

OpenStudy (anonymous):

so far

OpenStudy (anonymous):

Now, it doesn't matter wether you take the root of 8.16 and divide it by two or you can do it the other way around. This is one of the fundamental arithmetical laws of mathematics: The equasion can be rewritten to\[\sqrt{8.16} \times {1\over2}\]

OpenStudy (anonymous):

Still with me?

OpenStudy (anonymous):

lol, not sure if i am now

OpenStudy (anonymous):

Dividing by two is the same as multiplying by one half, never heard of that?

OpenStudy (anonymous):

yes that makes sense

OpenStudy (anonymous):

OK, that is the only thing I have done so far.

OpenStudy (anonymous):

Now, you can do whatever you want with one half, as long as it stays one half technicaly. For example you can make it 2/4 or 30/60, it doesn't matter to the outcome of the equation, are you following?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great! So why don't we say we don't say one half, we say \[{1\over \sqrt{4}} = \sqrt{1\over4} = {1\over2}\]] because that is the same as one half, right?

OpenStudy (anonymous):

Following?

OpenStudy (anonymous):

yes I believe so

OpenStudy (anonymous):

Ok, so now we rewrite the equation to \[\sqrt{8.16}\times \sqrt{1\over4}\] Ok?

OpenStudy (anonymous):

Now, it is laws of mathematics we apply: Because under the laws of arithmetic if the operation is divison or multiplication, we can perform these tasks BEFORE we take their roots. So we can rewrite to: \[\sqrt{8.16\times{1\over4}} = \sqrt{8.16\over4}\]

OpenStudy (anonymous):

And there we go, this equals: \[\sqrt{2.04}\] which is as far as we can simplify without using a calculator.

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

so i could divide by 4 even if the number was 9.69?

OpenStudy (anonymous):

Yes of course, as you saw, the four comes solely from one half, it has not anything to do with the other number.

OpenStudy (anonymous):

But I can ask you this: if you were supposed to divide by three in the beginning, what would you have to divide by then under the root?

OpenStudy (anonymous):

my guess would be 6

OpenStudy (anonymous):

no, what does a root do? it finds the quadratic exponent of the number so you to get the same result as before you have to square the number. So three squared would be?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

yes! to take the divisor three under the root, you would have to square it, and thus divide by nine in the root.

OpenStudy (anonymous):

ok. i understand now, after you do that, what are the operations to complete the solution?

OpenStudy (anonymous):

Well, you simply do what the equation says, in the case above, you divide 8.16 by four, take its root and subtract it from 2.02

OpenStudy (anonymous):

if you are not allowed to use a calculator, simply write 2.02 ± root of (2.04)

OpenStudy (anonymous):

im allowed to use a calculator

OpenStudy (anonymous):

Well there you go, then take the root of 2.04 and add it to 2.02 for the first x, and subtract it to get the second x

OpenStudy (anonymous):

There are always two results for a quadratic equation.

OpenStudy (anonymous):

im not sure how to get the root with the calculator though, i a using windows 7 calculator, scientific view

OpenStudy (anonymous):

is it 1.42?

OpenStudy (anonymous):

puh, well i think you have to type in the number, and press the root symbol, did it work? should give: 1.42....

OpenStudy (anonymous):

yes, thats what the calculator said? well technically it is 4.3 because it's 1.428... so you have to round up to 1.43

OpenStudy (anonymous):

sorry, 1.43

OpenStudy (anonymous):

Alright, is everything clear so far? I have to leave now, hope I could help a little.

OpenStudy (anonymous):

it's giving me 1.418

OpenStudy (anonymous):

you have been abig help thank you

OpenStudy (anonymous):

no worries! take care